OpenStudy (anonymous):

what is 44% of 80 round to the nearest whole number what do I do.

OpenStudy (anonymous):

turns out to be 35.20

OpenStudy (anonymous):

80 * 0.44 = 35.2

OpenStudy (anonymous):

\[ 1\% = 80 / 100 = 0.8 \\ 4 \% = 1\% \cdot 4 = 0.8 \cdot 4 = 3.2 \\ 40 \% = 1\% \cdot 40 =(1\% \cdot 4) \cdot 10 = (3.2) \cdot 10 = 32\\ 40 \% + 4\% = 32 + 3.2 = 35.2 \]

OpenStudy (anonymous):

Is that the answer.

OpenStudy (anonymous):

35.2 is the answer, but I tried to show the way too hmmph

OpenStudy (anonymous):

It is all whole numbers on my assignment.

OpenStudy (anonymous):

I'm sure you could make it shorter. Just tried to show it the way I see it. you could calculate this in your head I guess.. not too complicated in this case. Anyway, if that's not what you wanted then I apologize.

OpenStudy (anonymous):

Thats ok, you are right.

OpenStudy (anonymous):

The answers say 35, 44, 32, and 40

OpenStudy (anonymous):

Do you a twitter.

OpenStudy (anonymous):

the nearest whole number of 32.5 from the list is 32, right? hehe

OpenStudy (anonymous):

thank you now I know what you meant by that.

OpenStudy (anonymous):

sure np, good luck with the rest =]

OpenStudy (anonymous):

what does np stand for.

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

oh. ok.

OpenStudy (anonymous):

Do you have a twiiter.

OpenStudy (anonymous):

no, why?

OpenStudy (anonymous):

oh becuase I was going add you so when ever I need help I can talk with you about it.

OpenStudy (anonymous):

I have skype if you want. or you could just send me msg on this website

OpenStudy (anonymous):

I don't have a skype but i have been trying to ask my brother to make me one but he won't.

OpenStudy (anonymous):

msn maybe?

OpenStudy (anonymous):

The question I am on now says After a meal costing $12.50 you decide to tip the waiter 15%. How much will you tip the waiter. Round to the nearest cent.

OpenStudy (anonymous):

What do u mn by that.

OpenStudy (anonymous):

reading

OpenStudy (anonymous):

reading what.

OpenStudy (anonymous):

Dam, some long responses.

OpenStudy (anonymous):

Ok.. \[ 100\% \div 5\% = 20 \\ 5 \% = 100\% \div 20 = 12.5 \div 20 = 6.25 \div 10 = 0.625 \\ 15\% = 5\% \times 3 = 0.625 \times 3 = \\ = (0.6 \times 3) + (0.02 \times 3) + (0.005 \times 3) = \\ = (1.8) + (0.06) + (0.015) = 1.875 \]

OpenStudy (anonymous):

I don't understand it really good I can but I can't.

OpenStudy (anonymous):

ok. let's simplify this. can you find \(1\%\) of 12.5?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

ok, what is it then? =]

OpenStudy (anonymous):

really I don't know wat that is. but was that the right answer you wanted me to answer.

OpenStudy (anonymous):

Do you know what a % means?

OpenStudy (anonymous):

yes I know wat % this means.

OpenStudy (anonymous):

I'm not trying to mock, i'm asking seriously.

OpenStudy (anonymous):

I am sorry for doing that to you.

OpenStudy (anonymous):

it's ok i don't mind. Let's see. We know the WHOLE sum is 12.5 right? so we say that's our 100%. the 1% is therefore \(\frac{1}{100}\) of it. right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

if we divide 12.5 by 100 what do you get then?

OpenStudy (anonymous):

sorry but do you have a facetime.

OpenStudy (anonymous):

No, sorry =( Do you get what I mean though? \[ 12.5 \div 100 = \;? \]

OpenStudy (anonymous):

you should get 0.125 correct.

OpenStudy (anonymous):

yes very good =]

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

Now, we need to multiply that by 15 hehe to get from 1% to 15%

OpenStudy (anonymous):

r u saying to multiply 15 by one.

OpenStudy (anonymous):

it's abit ugly, but the cleanest way I know to do this is to simply break this number to parts: \[ 0.125 = \\ \quad\quad\quad\quad 0.1 + \\ \quad\quad\quad\quad 0.02 + \\ \quad\quad\quad \quad 0.005 \]

OpenStudy (anonymous):

is that clear so far?

OpenStudy (anonymous):

ok add that. Right

OpenStudy (anonymous):

now, if we want to multiply our 1% (which is 0.125 remember!) then we do \[ 0.125 \times 15 = \\ \quad\quad\quad\quad 0.1 \times 15 + \\ \quad\quad\quad\quad 0.02 \times 15 + \\ \quad\quad\quad \quad 0.005 \times 15 \]

OpenStudy (anonymous):

Could you solve each one of those?

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

Is it 1.875

OpenStudy (anonymous):

yes =]

OpenStudy (anonymous):

thank you very much.

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

do I round that.

OpenStudy (anonymous):

are u a student or a teacher.

OpenStudy (anonymous):

Do they ask you to?

OpenStudy (anonymous):

well, neither.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

what are u than.

OpenStudy (anonymous):

Well, they say to 'nearest cent'. a cent is a $0.01 right? so \(1.875 \to 1.87 \)

OpenStudy (anonymous):

it could be \(1.875 \to 1.88\). it's actually more common to round 5 up..

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

U r correct on the spot great job.

OpenStudy (anonymous):

Well, those are kinda the easy stuff heh. Some of the questions here are.... way beyond my level.

OpenStudy (anonymous):

really and u be getting them right.

OpenStudy (anonymous):

and heh.

OpenStudy (anonymous):

So, could you try the rest on your own?

OpenStudy (anonymous):

I might be able to do them on my own I will try.

OpenStudy (anonymous):

ok good luck. I might be going to sleep soon, it's midnight in here.. If you get into troubles then send me PM and i'll help you tomorrow or something

OpenStudy (anonymous):

I will right u back when ever I need help ok?.......?????????????

OpenStudy (anonymous):

sure fine

OpenStudy (anonymous):

any questions.